OpenStudy (anonymous):

How do you find the value of D? 8-2d+4d=12

OpenStudy (anonymous):

first collect like terms 8 + 2d = 12 2d = 4 d = 2

OpenStudy (anonymous):

\[8-2d+4d=12\] \[8+2d=12\] \[2d=4\] \[d=2\]

OpenStudy (anonymous):

that taught me absolutly nothing. but thanks...

OpenStudy (anonymous):

what you don't understand here? -2+4d=2d correct?